Aristotle plans to save for a down payment on a house he will buy 8 years later. He will be able to invest 107,9 tuuay wram e explain please, with financial calculator
Answer #1

Amount Invested Today = $109040

Future Value(FV) formula = P*(1+-)nt

where, P= Principal amount Invested ;= $ 109040

r = Interest rate ;= 0.07

n = No of Period in a year ;= 4 (Compounded Quarterly)

t= Time period over years ; = 8

FV = 109040 + (1 + 0.07 4-5

= 109040 *(1.0175)

= 109040 + (1.742213)

= $ 189,870.90

Since, Aristotle requires $ 189,791 in 8 years & he is able to generate this much amount from investing $109040 at 7 % interest compounded quarterly for 8 years. Hence, Yes his target will be able to meet at the end of 8 years.